40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Sep 7 23:49:14 2020, gautam, Update, BHD, A first look at RF44 scheme 
    Reply  Tue Sep 8 01:31:43 2020, Koji, Update, BHD, A first look at RF44 scheme 
    Reply  Wed Sep 9 00:05:18 2020, gautam, Update, BHD, More notes on the RF44 scheme IMG_3397.JPGphaseNoisePSD.pdfmagnitudeHist.pdfLOpowerDrift.png
       Reply  Tue Sep 15 22:11:52 2020, gautam, Update, BHD, More notes on the RF44 scheme RF44.pdf
Message ID: 15562     Entry time: Mon Sep 7 23:49:14 2020     Reply to this: 15563   15565
Author: gautam 
Type: Update 
Category: BHD 
Subject: A first look at RF44 scheme 

Summary:

Over the last couple of days, I've been working towards getting the infrastructure ready to test out the scheme of sensing (and eventually, controlling) the homodyne phase using the so-called RF44 scheme. More details will be populated, just quick notes for now before I forget.

  • LO beam with RF sidebands needed to be re-coupled into collimator, it wasn't seated tightly and just touching the fiber completely destroyed the alignment.
  • HWP installed before said collimator - IMC wants s-polarized light whereas the IFO field is p-polarized.
  • After my work, the numbers were: ~1.47mW input to collimator, ~1.07mW out of collimator on AS table, ~1mW making it to the BHD board. All seem like reasonable numbers to me.
  • 44 MHz signal synthesis - for now, I use a Marconi (10 MHz synced to Rb clock), I think we could also use a mixer+SLP50 to mix 11 and 55 MHz signals (which are easily available at the LSC rack) to generate this. I looked at Wenzel quadruplers, the specs don't suggest a quadrupler will do much better.
  • CDS model was modified to accept the phase-tracker output as an error signal for the homodyne phase control servo. Compile and install went smooth but I opted against a model restart tonight, I'll do it tmrw.
  • Some trials were done with the Michelson locked to a dark fringe (as was done for the case of the DC LO field beating with the 55 MHz sideband). While the overall spectrum lines up fairly well with earlier trials, the signal looks somewhat more "discontinuous" in its traversal of I/Q space, and it never quite goes to 0. Some offset? What does this mean for locking? More investigations needed....
ELOG V3.1.3-